# Barriers that Programming Instructors Face While Performing Emergency Pedagogical Design to Shape Student-AI Interactions with Generative AI Tools

### v1

@@ Abstract.
copyright: acmlicensedjournalyear: 2026doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2026; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Human-centered computing Empirical studies in HCI
Figure 1. In contrast to user experience designers (left) who can directly observe users and modify interfaces to alter user behaviors, instructors (right) engage in emergency pedagogical design – they make attempts to improve student-AI interactions with limited visibility into student behaviors and lack of control over commercial GenAI interfaces. :
@@ 1. Introduction
Instructors around the world are reacting to the availability of highly capable generative artificial intelligence (GenAI) tools like ChatGPT (openaiIntroducing2022, 45), Claude (goldmanOpenAI2023, 22), and Gemini (pichaiIntroducing2023, 48), which have raised questions around the value of formative assessments like take-home essays and programming assignments that can easily be completed by these tools (walshEveryone2025, 68). Students already interact with GenAI via freely available GenAI applications, sometimes even without intending to do so because GenAI outputs are now included by default in web search results (reidGenerative2024, 54), word processing software (spataroIntroducing2023, 62), and code editors (hollandAnnouncing2024, 27). Since unrestricted use of GenAI tools can hinder student learning (bastaniGenerative2024a, 4, 76), instructors are now finding themselves in the unenviable position of trying to structure student interactions with GenAI tools that weren’t originally designed for instruction.
@@
Drawing an analogy to emergency remote teaching where instructors needed to abruptly offer courses online in response to the COVID-19 pandemic (hodgesdifference2020, 26), we characterize the work that instructors are now performing as emergency pedagogical design. In addition to all of the previous responsibilities of running a course – creating course materials, lecturing, grading, and managing course staff – instructors now suddenly need to encourage students to interact with GenAI tools in ways that help rather than hinder learning. For example, some instructors banned GenAI usage entirely (johnsonChatGPT2023, 29), while others found it important to deliberately demonstrate its limitations (rooseDont2023, 57). Although these two approaches are different, in both cases instructors made an intervention (e.g. banning GenAI) and then tried to evaluate the efficacy of these interventions on student-AI behaviors.
In this framing, instructors who are designing student-AI interactions share some similarities to user experience (UX) designers for software. For example, both instructors and UX designers establish goals for a target audience: instructors want to maximize student learning, while a designer might want to encourage users to engage with their application. To accomplish these goals, both groups observe how people interact with software and use this evidence to encourage user behavior that is more aligned with their goals. However, there are differences as well, as depicted in Figure 1. For example, UX designers are trained in design thinking and user research methods, while instructors primarily focus on pedagogy and their discipline-specific research area. But perhaps the most salient difference is that designers often have the ability to change the interface itself, while instructors lack control over most GenAI interfaces. Instead, instructors manipulate course policies, materials, assessments, and classroom culture in order to promote desirable student-AI interactions (kasneciChatGPT2023, 30, 10).
@@
This paper examines undergraduate computing instructors as a critical early case where GenAI’s impact arrived to observe emergency pedagogical design in practice. In particular, in this setting we consider computing instructors to be lead users (vonhippelLead1986, 67) – users who are trying to design solutions for their own experienced needs. For example, computing instructors are attempting to shape student-AI interactions by creating alternative interfaces that wrap off-the-shelf GenAI tools (e.g. (kazemitabaarCodeAid2024, 33, 39)), making them a potential source of need-forecasting for instructors in other domains. This leads to our central research question: How do computing instructors conduct emergency pedagogical design, and what barriers do they encounter?
@@
To address this question, we interviewed 13 computing instructors who had made substantial attempts to influence student-AI interactions by integrating GenAI into their course assignments, assessments, and infrastructure. To gauge broader patterns among a more diverse set of instructors, we also surveyed 169 computing instructors about their perspectives towards integrating GenAI tools into their courses. Instructors repeatedly encountered five barriers as they engaged in emergency pedagogical design: fragmented buy-in for revising courses; policy crosswinds from non-prescriptive institutional guidance; implementation challenges as instructors attempt interventions; assessment misfit as student-AI interactions are only partially visible to instructors; and lack of resources, including time, staffing, and paid tool access. This study was conducted in mid-2025, about two and a half years after the initial release of well-known GenAI tools like ChatGPT and Github Copilot. Thus, our work captures a rare moment in time where computing instructors have been able to make initial attempts to shape student-AI interactions, but approaches have not yet converged on a set of best practices. To spur the community to use this unique opportunity to influence how instructors approach student-AI interaction, we conclude this paper with a set of open research questions for HCI (Section 5.1) and recommendations for academic institutions and funders (Section 5.2).
@@ 2.1. GenAI Adoption and Policy Responses in Computing Education
The rapid development of GenAI models has created both opportunities and challenges for computing education. In particular, the launch of AI models that can generate code like GitHub Copilot (friedmanIntroducing2021, 21) in 2021 and ChatGPT in 2022 (openaiIntroducing2022, 45) sparked a flurry of research that found that these tools could solve problems across the undergraduate computing curricula, for example in introductory programming courses (finnie-ansleyRobots2022, 16, 53, 59, 69, 61, 58), data structures courses (finnie-ansleyMy2023, 17), programming competitions (liCompetitionlevel2022, 38), and software engineering courses (choudhuriHow2024, 8). Instructors reacted to this in the short-term by banning the use of GenAI tools in their course policies and reduced the weight of non-proctored assessments (lauBan2023, 36). As these tools became widely accessible, however, instructors increasingly acknowledged that simply changing course policies was not enough; the content and learning goals of their courses also needed to adapt. This change in perspective was motivated by observations that students used GenAI tools regardless of policy (adninExamining2025a, 2, 68, 11), that GenAI coding assistants had already become commonplace in industry (mehtaquarter2025, 44, 66), and that GenAI tools could provide new opportunities to improve programming pedagogy (pratherRobots2023, 51, 13, 20, 52).
Some instructors have begun to integrate GenAI directly into their courses. For instance, Vadaparty et al. redesigned an introductory programming course to incorporate GenAI from the start (vadapartyCS1LLM2024, 65), and Benario et al. did the same for a software engineering course (gorsonbenarioUnlocking2025, 23). However, efforts like this seem to be the exception, not the norm. In a survey of computing instructors in 2024, 75% of surveyed faculty believed that their courses’ learning objectives should change in response to GenAI, yet only 35% reported actually incorporating GenAI into their courses (pratherHype2025, 50). A similar pattern appears outside computer science: a global survey conducted in early 2025 reported that 88% of faculty use AI sparingly or not at all, even though 86% expect to use AI for teaching in the future (digitaleducationcouncilWhat2025, 10). Reported reasons for this gap include distrust of GenAI (borelaWhat2025, 5, 41, 7, 47), limited understanding of its capabilities (tanMore2024, 63), concerns about effects on learning outcomes (zastudilGenerative2023, 75, 19, 25), and concern for potential student overreliance (sheardInstructor2024, 60, 24, 14, 10).
@@ 2.2. Systems that Support Learning to Program with GenAI
Since GenAI use for programming has become more common in the past few years, research on how novices learn programming with GenAI is in its early stages with both positive and negative results. For instance, Kazemitabaar et al. found that using a GenAI coding assistant can improve novices’ ability to write code even on later tasks without access to GenAI (kazemitabaarStudying2023, 31). However, GenAI tools impose additional metacognitive requirements on users (tankelevitchMetacognitive2024, 64), which can reinforce learning challenges for struggling students in the context of programming (pratherWidening2024, 49).
HCI research has proposed system designs to mitigate the potential harms of GenAI coding tools while maximizing their benefits for programming learners. For example, GenAI coding tools are often designed to quickly produce working code (jimenezSWEbench2024, 28), not to help learners understand how programs work. In response, researchers have proposed GenAI tools with guardrails that offer conceptual help rather than writing code (kazemitabaarCodeAid2024, 33, 39), and that ask students to “teach” a chatbot (rogersPlaying2025a, 56). To improve user understanding of code, Ivie adds lightweight explanations of generated code (yanIvie2024, 71), Leap surfaces live runtime values (ferdowsiValidating2024a, 15), and WaitGPT visualizes the flow of data through a program (xieWaitGPT2024, 70).
Novices can struggle with writing effective prompts (zamfirescu-pereiraWhy2023, 74) and often omit important detail in their prompts when trying to generate code (lucchettiSubstance2024, 40). In contrast, systems that guide users to decompose problems can increase perceived control and ease of use (kazemitabaarImproving2024, 32). Along this line of work, DBox guides learners to decompose their algorithms before writing code(maDBox2025a, 42), CoLadder provides an interface to manipulate a tree of prompts and generated code (yenCoLadder2024, 73), iGPT iteratively asks followup questions to learners to produce more complete program generation plans (yehBridging2025, 72), and Prompt Problems automatically grade learner prompts by checking generated code against test cases (dennyPrompt2024, 12). Some systems even remove the need for prompting altogether by offering proactive suggestions (chenNeed2025, 6, 34) or by offering a simplified interface to generate hints (roestNextStep2024, 55).
@@ 3. Methodology
To understand the characteristics of emergency pedagogical design, we conducted semi-structured interviews with 13 full-time instructors teaching undergraduate-level computing courses. We were specifically interested in understanding prior experiences of designing student-AI interactions rather than merely perceptions of possible interventions. To accomplish this, we selected interviewees who self-reported that they had made changes to their course materials because of student-AI usage, and had also taught these updated materials to students already. We filtered out instructors who had not yet made changes to their course or had only made changes to their course policies, since we felt this group of instructors had already been sufficiently studied in prior work (e.g. (lauBan2023, 36)). Interviewees were recruited via professional networks, email lists for computing instructors, and personal contacts, using snowball and purposive sampling (palinkasPurposeful2015, 46) that was iterative until saturation. The interviews were conducted in May and June 2025. All interviews were held in English and participants were not financially compensated for participating in the study.
@@ 3.2. Rationale for Our Interview Protocol
Our protocol design was based on several theoretical considerations. Since we were specifically interested in what instructors actually did (in contrast to what they wanted to do), we drew from critical incident technique (flanagancritical1954, 18), a qualitative method for collecting and analyzing events that had substantial impact on outcomes. In our context, this means that we spent the majority of the interview asking instructors to reflect on the most memorable interventions that they performed in their courses, with the expectation that talking about vivid experiences would encourage interviewees to share more relevant details. Note that we did not define “memorable” as “positive”; we were equally interested in unsuccessful experiences attempting to design student-AI interactions and made this clear to our interviewees as part of our protocol.
We also grounded our conversations around concrete pieces of course materials, such as take-home assignments or exam questions. This part of the protocol was inspired by the cognitive walkthrough (mahatodyState2010, 43) methodology in HCI, which recommends discussing a shared artifact to help elicit greater quality and quantity of observations. To reduce the risk of interviewees fixating on minute details of their course materials, we also asked interviewees to make higher-level reflections after each line of questioning.
Lastly, when designing individual interview questions, we applied best practices to reduce bias, such as using neutral wording and avoiding jargon (adamsConducting2015, 1). In our protocol, we also did not mention specific AI tools to avoid priming or anchoring biases, but if participants brought up a specific tool of their own accord, we asked followup questions about that tool.
@@ 3.3. Interview Participant Backgrounds
Table 1. We interviewed 13 computing instructors to understand their experiences conducting emergency pedagogical design for their courses. Gender = Male (M), Female (F), Non-binary (NB). Years = number of years of experience as a full-time instructor.
@@
P01 F US 12 Public PhD-granting Algorithms
P02 M US 26 Private undergrad-only Intro to programming
P03 M Canada 5 Public PhD-granting Data structures
P04 M US 6 Public PhD-granting Intro to programming, Software engineering
P05 M US 23 Public PhD-granting Intro to programming
P06 NB US 11 Private PhD-granting Intro to programming
P07 M US 28 Private PhD-granting Software design
P08 F US 9 Public PhD-granting Intro to programming
P09 M New Zealand 12 Public PhD-granting Intro to programming
P10 F US 22 Public PhD-granting Software engineering
P11 F US 9 Public PhD-granting Intro to programming
P12 M US 14 Public PhD-granting Web development
P13 M UK 42 Public PhD-granting Cybersecurity
Our interviewee backgrounds are summarized in Table 1. We recruited 13 participants (5 non-male). Most of our participants taught in the United States (10/13), while others taught in Canada (1/13), New Zealand (1/13), and the United Kingdom (1/13). Participants were all full-time faculty that taught undergraduate courses across the computing curriculum, including introductory programming, structures, algorithms, software engineering, and cybersecurity. There was considerable range in the number of years of full-time teaching experience (5–42 years, $\mu$=17 years). Participants primarily taught at public PhD-granting universities (10/13), while some taught at private PhD-granting universities (2/13) and undergraduate-only institutions (1/13).

### v2

@@ Abstract.
journalyear: 2026copyright: ccconference: Proceedings of the 2026 CHI Conference on Human Factors in Computing Systems; April 13–17, 2026; Barcelona, Spainbooktitle: Proceedings of the 2026 CHI Conference on Human Factors in Computing Systems (CHI ’26), April 13–17, 2026, Barcelona, Spaindoi: 10.1145/3772318.3790682isbn: 979-8-4007-2278-3/2026/04ccs: Human-centered computing Empirical studies in HCI
@@ 1. Introduction
Instructors around the world are reacting to the availability of highly capable generative artificial intelligence (GenAI) tools like ChatGPT (OpenAI, 2022), Claude (Goldman, 2023), and Gemini (Pichai and Hassabis, 2023), which have raised questions around the value of formative assessments like take-home essays and programming assignments that can easily be completed by these tools (Walsh, 2025). Students already interact with GenAI via freely available GenAI applications, sometimes even without intending to do so because GenAI outputs are now included by default in web search results (Reid, 2024), word processing software (Spataro, 2023), and code editors (Holland, 2024). Since unrestricted use of GenAI tools can hinder student learning (Bastani et al., 2024; Zhai et al., 2024), instructors are now finding themselves in the unenviable position of trying to structure student interactions with GenAI tools that weren’t originally designed for instruction.
@@
Drawing an analogy to emergency remote teaching where instructors needed to abruptly offer courses online in response to the COVID-19 pandemic (Hodges et al., 2020), we characterize the work that instructors are now performing as emergency pedagogical design. In addition to all of the previous responsibilities of running a course – creating course materials, lecturing, grading, and managing course staff – instructors now suddenly need to encourage students to interact with GenAI tools in ways that help rather than hinder learning. For example, some instructors have banned GenAI usage entirely (Johnson, 2023), while others have found it valuable to deliberately demonstrate its limitations (Roose, 2023). One clear similarity between emergency remote teaching and emergency pedagogical design is that both arose in response to a disruptive event: in the former, a global pandemic, and in the latter, the creation of free-to-use, highly capable GenAI tools. One notable difference is that emergency remote teaching was generally considered to be a temporary shift in teaching modality until the COVID-19 virus was under control, but there is general consensus that GenAI tools will have a role in education even in the long-term (Baidoo-Anu and Owusu Ansah, 2023). Thus, by studying instructors’ current responses to GenAI tools, we can deepen understanding of the practices that may soon be normalized in everyday instruction.
@@
This paper examines undergraduate computing instructors as a critical early case where GenAI’s impact arrived to observe emergency pedagogical design in practice. In particular, in this setting we consider computing instructors to be lead users (Von Hippel, 1986) – users who are trying to design solutions for their own experienced needs. For example, computing instructors are attempting to shape student-AI interactions by creating alternative interfaces that wrap off-the-shelf GenAI tools (e.g. (Kazemitabaar et al., 2024b; Liffiton et al., 2023)), making them a potential source of need-forecasting for instructors in other domains. This leads to our central research question: How do computing instructors conduct emergency pedagogical design, and what barriers do they encounter?
@@
To address this question, we interviewed 13 computing instructors who had made substantial attempts to influence student-AI interactions by integrating GenAI into their course assignments, assessments, and infrastructure. To gauge broader patterns among a more diverse set of computing instructors, including instructors at Minority-Serving Institutions (MSIs) and Historically Black Colleges and Universities (HBCUs), we also surveyed 169 instructors about their perspectives towards integrating GenAI tools into their courses. These instructors repeatedly encountered five barriers as they engaged in emergency pedagogical design: fragmented buy-in for revising courses; policy crosswinds from non-prescriptive institutional guidance; implementation challenges as instructors attempt interventions; assessment misfit as student-AI interactions are only partially visible to instructors; and lack of resources, including time, staffing, and paid tool access. This study was conducted in mid-2025, about two and a half years after the initial release of well-known GenAI tools like ChatGPT and GitHub Copilot. Thus, our work captures a rare moment in time where computing instructors have been able to make initial attempts to shape student-AI interactions, but approaches have not yet converged on a set of best practices. To spur the community to use this unique opportunity to influence how instructors approach student-AI interaction, we conclude this paper with a set of open research questions for HCI (Section 5.1), recommendations for academic institutions and funders (Section 5.2), and a reflection on lessons learned from emergency remote teaching that may also apply in this context (Section 5.3).
@@ 2.1. GenAI Adoption and Policy Responses in Computing Education
The rapid development of GenAI models has created both opportunities and challenges for computing education. In particular, the launch of AI models that can generate code like GitHub Copilot (Friedman, 2021) in 2021 and ChatGPT in 2022 (OpenAI, 2022) sparked a flurry of research that found that these tools could solve problems across the undergraduate computing curricula, for example in introductory programming courses (Finnie-Ansley et al., 2022; Reeves et al., 2023; Savelka et al., 2023b; Wermelinger, 2023; Singla, 2023; Savelka et al., 2023a), data structures courses (Finnie-Ansley et al., 2023), programming competitions (Li et al., 2022), and software engineering courses (Choudhuri et al., 2024). Instructors reacted to this in the short-term by banning the use of GenAI tools in their course policies and reduced the weight of non-proctored assessments (Lau and Guo, 2023). As these tools became widely accessible, however, instructors increasingly acknowledged that simply changing course policies was not enough; the content and learning goals of their courses also needed to adapt. This change in perspective was motivated by observations that students used GenAI tools regardless of policy (Adnin et al., 2025; Walsh, 2025; Cu and Hochman, 2023), that GenAI coding assistants had already become commonplace in industry (Mehta, 2025; Vanian, 2025), and that GenAI tools could provide new opportunities to improve programming pedagogy (Prather et al., 2023; Denny et al., 2024b; Franklin et al., 2025; Raihan et al., 2025).
Some instructors have begun to integrate GenAI directly into their courses. For instance, Vadaparty et al. redesigned an introductory programming course to incorporate GenAI from the start (Vadaparty et al., 2024), and Benario et al. did the same for a software engineering course (Gorson Benario et al., 2025). However, efforts like this seem to be the exception, not the norm. In a survey of computing instructors in 2024, 75% of surveyed faculty believed that their courses’ learning objectives should change in response to GenAI, yet only 35% reported actually incorporating GenAI into their courses (Prather et al., 2025). A similar pattern appears outside computer science: a global survey conducted in early 2025 reported that 88% of faculty use AI sparingly or not at all, even though 86% expect to use AI for teaching in the future (Council, 2025). Reported reasons for this gap include distrust of GenAI (Borela et al., 2025; Lyu et al., 2025; Choi et al., 2023; Park and Ahn, 2024), limited understanding of its capabilities (Tan and Subramonyam, 2024), concerns about effects on learning outcomes (Zastudil et al., 2023; Foroughi et al., 2024; Harvey et al., 2025), and concern for potential student overreliance (Sheard et al., 2024; Haque and Hundhausen, 2025; ElSayary, 2024; Council, 2025).
@@ 2.2. Systems that Support Learning to Program with GenAI
Since GenAI use for programming has become more common in the past few years, research on how novices learn programming with GenAI is in its early stages with both positive and negative results. For instance, Kazemitabaar et al. found that using a GenAI coding assistant can improve novices’ ability to write code even on later tasks without access to GenAI (Kazemitabaar et al., 2023). However, GenAI tools impose additional metacognitive requirements on users (Tankelevitch et al., 2024), which can reinforce learning challenges for struggling students in the context of programming (Prather et al., 2024).
HCI research has proposed system designs to mitigate the potential harms of GenAI coding tools while maximizing their benefits for programming learners. For example, GenAI coding tools are often designed to quickly produce working code (Jimenez et al., 2024), not to help learners understand how programs work. In response, researchers have proposed GenAI tools with guardrails that offer conceptual help rather than writing code (Kazemitabaar et al., 2024b; Liffiton et al., 2023), and that ask students to “teach” a chatbot (Rogers et al., 2025). To improve user understanding of code, Ivie adds lightweight explanations of generated code (Yan et al., 2024), Leap surfaces live runtime values (Ferdowsi et al., 2024), and WaitGPT visualizes the flow of data through a program (Xie et al., 2024).
Novices can struggle with writing effective prompts (Zamfirescu-Pereira et al., 2023) and often omit important detail in their prompts when trying to generate code (Lucchetti et al., 2024). In contrast, systems that guide users to decompose problems can increase perceived control and ease of use (Kazemitabaar et al., 2024a). Along this line of work, DBox guides learners to decompose their algorithms before writing code(Ma et al., 2025), CoLadder provides an interface to manipulate a tree of prompts and generated code (Yen et al., 2024), iGPT iteratively asks followup questions to learners to produce more complete program generation plans (Yeh et al., 2025), and Prompt Problems automatically grade learner prompts by checking generated code against test cases (Denny et al., 2024a). Some systems even remove the need for prompting altogether by offering proactive suggestions (Chen et al., 2025; Khurana et al., 2025) or by offering a simplified interface to generate hints (Roest et al., 2024).
@@ 3. Methodology
To understand the characteristics of emergency pedagogical design, we conducted semi-structured interviews with 13 full-time instructors teaching undergraduate-level computing courses. We were specifically interested in understanding prior experiences of designing student-AI interactions rather than merely perceptions of possible interventions. To accomplish this, we selected interviewees who self-reported that they had made changes to their course materials because of student-AI usage, and had also taught these updated materials to students already. We filtered out instructors who had not yet made changes to their course or had only made changes to their course policies, since we felt this group of instructors had already been sufficiently studied in prior work (e.g. (Lau and Guo, 2023)). Interviewees were recruited via professional networks, email lists for computing instructors, and personal contacts, using snowball and purposive sampling (Palinkas et al., 2015) that was iterative until saturation. The interviews were conducted in May and June 2025. All interviews were held in English and participants were not financially compensated for participating in the study.
@@ 3.2. Rationale for Our Interview Protocol
Our protocol design was based on several theoretical considerations. Since we were specifically interested in what instructors actually did (in contrast to what they wanted to do), we drew from critical incident technique (Flanagan, 1954), a qualitative method for collecting and analyzing events that had substantial impact on outcomes. In our context, this means that we spent the majority of the interview asking instructors to reflect on the most memorable interventions that they performed in their courses, with the expectation that talking about vivid experiences would encourage interviewees to share more relevant details. Note that we did not define “memorable” as “positive”; we were equally interested in unsuccessful experiences attempting to design student-AI interactions and made this clear to our interviewees as part of our protocol.
We also grounded our conversations around concrete pieces of course materials, such as take-home assignments or exam questions. This part of the protocol was inspired by the cognitive walkthrough (Mahatody et al., 2010) methodology in HCI, which recommends discussing a shared artifact to help elicit greater quality and quantity of observations. To reduce the risk of interviewees fixating on minute details of their course materials, we also asked interviewees to make higher-level reflections after each line of questioning.
Lastly, when designing individual interview questions, we applied best practices to reduce bias, such as using neutral wording and avoiding jargon (Adams, 2015). In our protocol, we also did not mention specific AI tools to avoid priming or anchoring biases, but if participants brought up a specific tool of their own accord, we asked followup questions about that tool.
@@ 3.3. Interview Participant Backgrounds
Table 1. We interviewed 13 computing instructors to understand their experiences conducting emergency pedagogical design for their courses. Gender = Male (M), Female (F), Non-binary (NB). Years = number of years of experience as a full-time instructor. Each course listed in the “Courses discussed” column is marked with (L) to denote a lower-division course or (U) for an upper-division course.
@@
P01 F US 12 Public PhD-granting Algorithms (U)
P02 M US 26 Private undergrad-only Intro to programming (L)
P03 M Canada 5 Public PhD-granting Data structures (L)
P04 M US 6 Public PhD-granting Intro to programming (L), Software eng. (U)
P05 M US 23 Public PhD-granting Intro to programming (L)
P06 NB US 11 Private PhD-granting Intro to programming (L)
P07 M US 28 Private PhD-granting Software design (U)
P08 F US 9 Public PhD-granting Intro to programming (L)
P09 M New Zealand 12 Public PhD-granting Intro to programming (L)
P10 F US 22 Public PhD-granting Software engineering (U)
P11 F US 9 Public PhD-granting Intro to programming (L)
P12 M US 14 Public PhD-granting Web development (U)
P13 M UK 42 Public PhD-granting Cybersecurity (L)
Our interviewee backgrounds are summarized in Table 1. We recruited 13 participants (5 non-male). Most of our participants taught in the United States (10/13), while others taught in Canada (1/13), New Zealand (1/13), and the United Kingdom (1/13). All participants held PhD degrees and were either tenure-line faculty (8/13) or nontenure-line faculty (5/13). With one exception that taught in an engineering department, all interviewees taught in computer science departments. Interviewees taught undergraduate courses at both introductory and upper levels, including introductory programming, data structures, algorithms, software engineering, and cybersecurity. There was considerable range in the number of years of full-time teaching experience (5–42 years, $\mu$=17 years). Participants primarily taught at public PhD-granting universities (10/13), while some taught at private PhD-granting universities (2/13) and undergraduate-only institutions (1/13).